A girl delivering newspaper covers her route by traveling 3.00 blocks west, 4.00 blocks north, and then 6.00 blocks east. (a) What is her resultant displacement? (b) What is the total distance she travels?
1
Expert's answer
2020-10-13T09:37:03-0400

a) To find her resultant displacement, it would be useful to create a brief illustration:



The displacement in this case is equal to the hypotenuse of a right triangle with sides equal to 3 (6 - 3) and 4 blocks. According to Pythagorean theorem,

"d=\sqrt{(3.00)^2+(4.00)^2}=5.00" blocks.


b) The total distance is just the sum of all distances regardless of the direction:

"s=3.00+4.00+6.00=13.0" blocks.


Answer: a) 5.00 blocks; b) 13.0 blocks